THE LOW-LYING BENDING VIBRONIC BANDS OF THE MgNC Ã^{2}Π – X̃ ^{2}Σ^{+} TRANSITION: ANALYSIS OF THE RENNER-TELLER VIBRONIC STRUCTURE

\underline{MASARU FUKUSHIMA} and TAKASHI ISHIWATA, Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

We have generated MgNC in supersonic free jet expansions, and measured the laser induced fluorescence excitation spectra of the Mg-N-C bending vibronic bands of the Ã ^{2}Π – X̃ ^{2}Σ^{+} transition. We analyzed 2_{0}^{1}, κ ^{2}Σ^{(+)} – and 2_{0}^{2}, κ ^{2}Π – ^{2}Σ^{+} bands and 2_{0}^{2}, μ ^{2}Π_{\frac{1}{2}} – ^{2}Σ^{+} sub-band. Through the molecular constants of the Ã (010) κ ^{2}Σ^{(+)}, (020) κ ^{2}Π and μ ^{2}Π_{\frac{1}{2}}

levels, the Renner-Teller vibronic structure on the ν_{2} bending mode in the Ã ^{2}Π state has been characterized. To analyze the vibronic structure, we need to consider the anharmonic term.